DETERMINATION OF THE ν_{9} BAND CENTER THROUGH THE FIRST HIGH-RESOLUTION ANALYSES OF THE ν_{8} and ν_{8} + ν_{9} SPECTRAL REGIONS OF ^{35}ClONO_{2}

\underline{J.-M. FLAUD}, J. ORPHAL, Laboratoire de Photophysique Moléculaire, C.N.R.S., Bât. 350, Université Paris-Sud, 91405 Orsay Cedex, France; W. J. LAFFERTY, Optical Technology Division, NIST, Gaithersburg, MD 20899, U.S.A; M. BIRK, and G. WAGNER, Institute for Opto-electronics, DLR, 82234 Oberpfaffenhofen, Germany.

A ro-vibrational analysis of the C-type ν_{8} fundamental band of ^{35}ClONO_{2} as well as the hot band, ν_{8} + ν_{9} − ν_{9}, has been carried out using a Fourier-transform infrared spectrum of natural chlorine nitrate. This spectrum was recorded with a resolution of 0.00094 cm^{−1} at a temperature of 190 K. Accurate upper state constants have been determined for both bands including the following band centers: ν_{0}(ν_{8}) = 711.20763(9) and ν_{0}(ν_{8} + ν_{9} − ν_{9}) = 714.9050(12) cm^{−1}. These constants have been used, together with a transition moment operator which takes into account the observed Herman-Wallis effect, to successfully model the experimental spectrum. Moreover the constants of the 8^{1}9^{1} state have been used to model the ν_{8} + ν_{9} band, allowing us to assign nearly 100 Q-branch transitions and, therefore, to determine the band center of the ν_{8} + ν_{9} band , ν_{0}(ν_{8} + ν_{9}) = 838.6269(15) cm^{−1}. Consequently we are able for the first time to obtain a precise band center of ν_{9} : ν_{0}(ν_{9}) = 123.7219(20) cm^{−1}. Finally this establishes unambiguously that the band at 714.9 cm^{−1} previously attributed to ν_{5} + ν_{7} − ν_{9} is actually the ν_{8} + ν_{9} − ν_{9} hot band.